# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Appendix E Reinforcement Learning Generated Subgraph Details
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $\mathbf{0.5293}\scriptstyle\pm\mathbf{0.0394}$ $0.5133\scriptstyle\pm 0.0379$ $0.5211\scriptstyle\pm 0.0379$ $0.3530\scriptstyle\pm 0.0344$ $0.7333\scriptstyle\pm 0.1155$ $0.7000\scriptstyle\pm 0.1000$
L3-FT(QA) + KG $0.4672\scriptstyle\pm 0.0472$ $0.4533\scriptstyle\pm 0.0231$ $0.4600\scriptstyle\pm 0.0346$ $0.2991\scriptstyle\pm 0.0296$ $0.8000\scriptstyle\pm 0.2000$ $0.8000\scriptstyle\pm 0.1000$
G-Retriever $0.4517\scriptstyle\pm 0.0318$ $0.4000\scriptstyle\pm 0.0173$ $0.4242\scriptstyle\pm 0.0237$ $0.2694\scriptstyle\pm 0.0193$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0000}$
@@
Table 35: Model performances on PRAD
@@
L3 $0111\scriptstyle\pm 0.0192 0.0$$033\scriptstyle\pm 0.0058 0.0$$051\scriptstyle\pm 0.0089 0.0$$26\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0145\scriptstyle\pm 0.0251 0.0$$033\scriptstyle\pm 0.0058 0.0$$054\scriptstyle\pm 0.0094 0.0$$27\scriptstyle\pm 0.0047 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0176\scriptstyle\pm 0.0261 0.0$$067\scriptstyle\pm 0.0058 0.0$$078\scriptstyle\pm 0.0083 0.0$$39\scriptstyle\pm 0.0042 0.000$$00\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(QA) $0.6021\scriptstyle\pm 0.0558$ $0.5600\scriptstyle\pm 0.0400$ $0.5796\scriptstyle\pm 0.0401$ $0.4088\scriptstyle\pm 0.0392$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9333\scriptstyle\pm 0.1155$
L3-FT(QA) + KG $\mathbf{0.6032}\scriptstyle\pm\mathbf{0.0191}$ $\mathbf{0.5867}\scriptstyle\pm\mathbf{0.0115}$ $\mathbf{0.5947}\scriptstyle\pm\mathbf{0.0102}$ $\mathbf{0.4232}\scriptstyle\pm\mathbf{0.0103}$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.1000$
G-Retriever $0.4749\scriptstyle\pm 0.0041$ $0.4433\scriptstyle\pm 0.0231$ $0.4584\scriptstyle\pm 0.0141$ $0.2974\scriptstyle\pm 0.0120$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$

### v2

@@ E.2 Human evaludation and LLM as judge
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $\mathbf{0.5293}\scriptstyle\pm\mathbf{0.0394}$ $0.5133\scriptstyle\pm 0.0379$ $0.5211\scriptstyle\pm 0.0379$ $0.3530\scriptstyle\pm 0.0344$ $0.7333\scriptstyle\pm 0.1155$ $0.7000\scriptstyle\pm 0.1000$
L3-FT(CRISPR) + Omics + KG $0.4672\scriptstyle\pm 0.0472$ $0.4533\scriptstyle\pm 0.0231$ $0.4600\scriptstyle\pm 0.0346$ $0.2991\scriptstyle\pm 0.0296$ $0.8000\scriptstyle\pm 0.2000$ $0.8000\scriptstyle\pm 0.1000$
G-Retriever + pre-GAT $0.4517\scriptstyle\pm 0.0318$ $0.4000\scriptstyle\pm 0.0173$ $0.4242\scriptstyle\pm 0.0237$ $0.2694\scriptstyle\pm 0.0193$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0000}$
RoG $0.4508\scriptstyle\pm 0.0506$ $0.4233\scriptstyle\pm 0.0982$ $0.4355\scriptstyle\pm 0.0772$ $0.2803\scriptstyle\pm 0.0614$ $0.6667\scriptstyle\pm 0.5774$ $0.4667\scriptstyle\pm 0.4041$
SubgraphRAG $0.4800\scriptstyle\pm 0.0000$ $0.4800\scriptstyle\pm 0.0000$ $0.4800\scriptstyle\pm 0.0000$ $0.3158\scriptstyle\pm 0.0000$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.7000\scriptstyle\pm 0.0000$
GNN-RAG $0.4777\scriptstyle\pm 0.0739$ $0.4433\scriptstyle\pm 0.1159$ $0.4589\scriptstyle\pm 0.0975$ $0.3011\scriptstyle\pm 0.0795$ $0.6000\scriptstyle\pm 0.5292$ $0.4667\scriptstyle\pm 0.4163$
@@
GALAX (Qwen2.5-7B) $0.4848\scriptstyle\pm 0.0391$ $0.5033\scriptstyle\pm 0.0231$ $0.4938\scriptstyle\pm 0.0313$ $0.3282\scriptstyle\pm 0.0280$ $0.9333\scriptstyle\pm 0.1155$ $0.8333\scriptstyle\pm 0.1155$
Table 43: Model performance on PRAD
@@
L3 + Omics $0111\scriptstyle\pm 0.0192 0.0$$033\scriptstyle\pm 0.0058 0.0$$051\scriptstyle\pm 0.0089 0.0$$26\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0145\scriptstyle\pm 0.0251 0.0$$033\scriptstyle\pm 0.0058 0.0$$054\scriptstyle\pm 0.0094 0.0$$27\scriptstyle\pm 0.0047 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0176\scriptstyle\pm 0.0261 0.0$$067\scriptstyle\pm 0.0058 0.0$$078\scriptstyle\pm 0.0083 0.0$$39\scriptstyle\pm 0.0042 0.000$$00\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics $0.6021\scriptstyle\pm 0.0558$ $0.5600\scriptstyle\pm 0.0400$ $0.5796\scriptstyle\pm 0.0401$ $0.4088\scriptstyle\pm 0.0392$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9333\scriptstyle\pm 0.1155$
L3-FT(CRISPR) + Omics + KG $0.6032\scriptstyle\pm 0.0191$ $0.5867\scriptstyle\pm 0.0115$ $0.5947\scriptstyle\pm 0.0102$ $0.4232\scriptstyle\pm 0.0103$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.1000$
G-Retriever + pre-GAT $0.4749\scriptstyle\pm 0.0041$ $0.4433\scriptstyle\pm 0.0231$ $0.4584\scriptstyle\pm 0.0141$ $0.2974\scriptstyle\pm 0.0120$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
RoG $0.5833\scriptstyle\pm 0.0404$ $0.5833\scriptstyle\pm 0.0404$ $0.5833\scriptstyle\pm 0.0404$ $0.4125\scriptstyle\pm 0.0410$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9333\scriptstyle\pm 0.0577$
SubgraphRAG $0.5931\scriptstyle\pm 0.1258$ $0.4500\scriptstyle\pm 0.0964$ $0.5042\scriptstyle\pm 0.0691$ $0.3389\scriptstyle\pm 0.0607$ $0.6667\scriptstyle\pm 0.5774$ $0.6333\scriptstyle\pm 0.5508$
GNN-RAG $0.6140\scriptstyle\pm 0.0138$ $0.6100\scriptstyle\pm 0.0173$ $0.6120\scriptstyle\pm 0.0156$ $0.4411\scriptstyle\pm 0.0163$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
@@
GALAX (Qwen2.5-7B) $\mathbf{0.6195}\scriptstyle\pm\mathbf{0.0292}$ $\mathbf{0.6133}\scriptstyle\pm\mathbf{0.0289}$ $\mathbf{0.6164}\scriptstyle\pm\mathbf{0.0290}$ $\mathbf{0.4459}\scriptstyle\pm\mathbf{0.0300}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$